OpenStudy (anonymous):

clear parentheis and comebine 2 (s+6)-(8s-t)+6t

OpenStudy (anonymous):

First multiply out the brackets 2(s+6) = 2s+12 -(8s-t) = -8s+t (Don't forget to distribute the - so that t becomes -t) +6t Bringing it all together: 2(s+6)-(8s-t) +6t = 2s+12- 8s+t +6t Grouping like terms (combine s terms together, likewise t terms, likewise the number term) gives: -6s+7t+12